Calculate the error in measured absorbance of a 1.3 micrometer thick sample with n= 2.7 and alpha= .2 per micronmeter, if the reflectance were neglected. Assume normal incidence.
Answer #1

Let us assume that the light is travelling from air (n1 = 1 to the slab n2 = 2.7)

Reflectance coefficient for normal incidence R = ((n1-n2)/(n1+n2))2

R = (1.7/3.7)2 =0.459 = 0.46

As we know that Transmittance +Absorbance +Reflectance = 1

Absorbance = 1 - (Transmittance +Reflectance )

If reflectance is neglected, Absorbance = 1 - (Transmittance)

so error in absobance = Reflectance.

Please see the below image on how we calculate the reflectanceDato R-Cue je.et斗叫ethin.-→o46 fa-absorption coe彬4 一ㄧ一 교 4ーーーフ AL Serban ce Ci-Re-ad nic A- A2 0-46 0-46 (6-s-052 20.S2 0 974

Hence, error in absorbance = 0.54